The probability that a hockey game will go into overtime is .3; Of the next 12 hockey games, what is the probability that more than 3 (P(x >3)) will go into overtime?
I think this is Poisson, but am not 100 percent sure. Would appreciate some guidance. Thank you.

Answer by stanbon(75887)

The probability that a hockey game will go into overtime is .3; Of the next 12 hockey games, what is the probability that more than 3 (P(x >3)) will go into overtime?
-----------------------
Binomial Problem with n = 12 and p(overtime) = 0.3
----
P(4<= x <=12) = 1 - P(0<= x <=3) = 1 - binomcdf(12,0.3,3)
---
= 1 - 0.4925
----
= 0.5075
=============
Cheers,
Stan H.
=============